undeadknight26 (undeadknight26):

Given the function f(x) = x3 + 2x2 - 3x - 5, what is the resulting function when f(x) is shifted to the right 1 unit?

undeadknight26 (undeadknight26):

f(x) + 1 = x3 + 2x2 - 3x - 4 f(x) − 1 = x3 + 2x2 - 3x - 6 f(x − 1) = x3- x2 - 4x - 1

undeadknight26 (undeadknight26):

f(x) + 1 = x3 + 2x2 - 3x - 4 f(x) - 1 = x3 + 2x2 - 3x - 6 f(x - 1) = x3- x2 - 4x - 1

undeadknight26 (undeadknight26):

I think A? @aaronq @agent0smith @ganeshie8 @iGreen

OpenStudy (anonymous):

f(0) = 1 (At x=0 the function is 1). I got this be substituting 0 for all the x's. So (0,1) is your first point. At x=1 the function is -4+4-2+1+1 = 0 So (1,0) is your second point. slope = (0-1) / (1-0) = -1 So the average rate of change for the function between x = 0 and x = 1 is -1

OpenStudy (igreen):

\( f(x) + 1 = x^3 + 2x^2 - 3x - 4\) is the same thing as the original

OpenStudy (igreen):

\( f(x) - 1 = x^3 + 2x^2 - 3x - 6 \) is the same as the original also..only the last option is different.

undeadknight26 (undeadknight26):

So the last option would be the correct answer...